### v2

# TabMGP: Martingale Posterior with TabPFN

Abstract
Bayesian inference provides principled uncertainty quantification but is often limited by challenges of prior and likelihood elicitation. The martingale posterior (MGP) (Fong et al., 2023) offers an alternative by replacing these requirements with a predictive rule. Additionally MGP focuses inference on parameters defined through a loss function. This framework is especially resonant in the era of foundation transformers; practitioners increasingly leverage models like TabPFN for their state-of-the-art capabilities, yet often require epistemic uncertainty for a scientific estimand $\theta$ that need not parameterise the model’s implicit latent model. The MGP provides the mechanism to recover these posterior distributions. We introduce TabMGP, an MGP built on TabPFN for tabular data. TabMGP produces credible sets with near-nominal coverage and often outperforms both handcrafted MGP constructions and standard Bayesian baselines.
1 Introduction
[…]
Figure 1: TabMGP for obtaining posterior samples of $\theta(F_{\infty})\mid z_{1:n}$, where each $z=(x,y)$ represents a covariate-response pair. We perform forward sampling with TabPFN to generate $L$ independent continuations of the observed dataset $z_{1:n}$. Since TabPFN does not model covariates, the forward sampling of $x_{i}$ is performed via a separate process, while TabPFN provides the conditional response $y_{i}\sim\mathrm{TabPFN}(\cdot\mid x_{i},z_{1:i-1})$. At the termination of each rollout $l$, we form the empirical measure $F_{N}^{(l)}=\frac{1}{N}\left(\sum_{i=1}^{n}\delta_{z_{i}}+\sum_{i=n+1}^{N}\delta_{z_{i}^{(l)}}\right)$ and collect $\theta(F_{N}^{(l)})$ as one approximate posterior sample.
Classical Bayesian inference provides a principled framework for uncertainty quantification, but its formulation hinges on the specification of a prior–likelihood pair. This requirement can be burdensome in practice. Priors are often placed on uninterpretable parameters, and “uninformative” defaults can dominate inference in data-scarce or high-dimensional regimes (Zhang et al., 2022). The problem is further compounded by the frequent misspecification of the likelihood in modern applications; even when a model is settled upon, approximating the resulting posterior—particularly in high dimensions—remains a formidable computational challenge.
These limitations have catalysed a wave of generalisations—ranging from Bayesian predictive inference (BPI) (Fortini and Petrone, 2025) to generalised Bayes (GB) (Bissiri et al., 2016) and optimisation-centric Bayes (see e.g. Knoblauch et al., 2022; Wild et al., 2023; Shen et al., 2025)—that collectively mark the emergence of a post-Bayesian paradigm. Within this paradigm, BPI shifts the focus from prior and likelihood elicitation to the sequence of predictions themselves. Specifically, the analyst provides a predictive rule, defined as a sequence of one-step-ahead predictive distributions $(P_{i})_{i\geq 0}$. BPI is post-Bayesian in the sense that the predictive rule need not correspond to a standard Bayesian posterior prediction distribution (PPD). Instead, the framework requires only that the sequence $(P_{i})_{i\geq 0}$ converges weakly to random limiting measure $F_{\infty}$ almost surely.
Then, we may ask for the posterior distribution of $F_{\infty}$ given data $z_{1:n}$ for some measurable set $A$. Here, $F_{\infty}$ represents the latent Bayesian model internalised by the predictive rule—a state of knowledge that remains implicit but can be queried through simulation. Specifically, we may obtain posterior samples of $F_{\infty}\mid z_{1:n}$ through a process of forward sampling: starting from the observed dataset $z_{1:n}$, one repeatedly draws the next observation from the predictive rule conditional on the data so far, append it to the dataset, and continue. In practice, the forward sampling process is terminated at some large $N$ such that $F_{N}$ serves as an approximate posterior sample of $F_{\infty}\mid z_{1:n}$.
Complementary to BPI is GB, a framework that argues we can update a prior to a posterior distribution over parameters that are connected to observations through a loss function rather than a likelihood. This allows for significant decision-theoretic flexibility: the object of interest $\theta$ is generally defined as the minimiser of an expected loss and need not index the underlying Bayesian model. While GB allows for decision-theoretic flexibility, it still requires the specification of an explicit prior.
The martingale posterior (MGP) emerges as a powerful synthesis of these directions, one that is especially resonant in the era of foundation transformers (Bommasani et al., 2022), which have recently begun to dominate tabular data tasks—a domain long considered the holdout of traditional machine learning, i.e., non-deep learning, methods. Practitioners confronted with a new tabular dataset will increasingly turn to pretrained models like TabPFN (Hollmann et al., 2022) rather than design a bespoke likelihood–prior pair and navigate the complexities of Markov chain Monte Carlo or variational inference. For these users, the parameters of scientific interest are effectively decoupled from the internals of the transformer, necessitating a framework that can recover the posterior distribution of a functional $\theta(F_{\infty})$ directly from the transformer’s predictive outputs. The MGP provides this mechanism: it treats the transformer as a predictive engine for the latent model $F_{\infty}$, while the GB loss function specifies the functional $\theta$ to be extracted.
The forward sampling mechanism we described above for $F_{\infty}\mid z_{1:n}$ provides the computational means for this recovery. The procedure remains the same except, at termination, we report $\theta(F_{N})$ as an approximate posterior sample from $\theta(F_{\infty})\mid z_{1:n}$. Note the sampling is well suited to transformer architectures: next-token prediction is the predictive rule, and forward sampling mirrors autoregressive generation, where each token is sampled from the conditional distribution given the context, appended to the context, and the process is repeated. This procedure is visualised in Figure 1.
Yet, despite this alignment, the martingale posterior literature has focused almost exclusively on handcrafted predictive rules that satisfy exact martingale properties (e.g., Fong et al., 2023; Ghalebikesabi et al., 2023; Huk et al., 2024); see Section 4 for a review. These approaches are theoretically rigorous, but they require problem-specific design choices to tune the predictive rule for the dataset at hand. Moreover, the emphasis on exact martingale structure overstates the importance of this condition: while the martingale property is a sufficient condition to guarantee the existence of a limiting law $F_{\infty}$, it is not necessary for delivering epistemic uncertainty for $\theta(F_{\infty})$. We argue that the reliance on this restrictive sufficient condition has acted as impediment to the growing field of MGP, preventing the application of MGPs to high-capacity models.
This motivates our central question: can modern foundation transformers be used as the predictive rule in martingale posteriors? We answer in the affirmative by introducing TabMGP, the first martingale posterior powered by a large pretrained foundation transformer. Specifically, we instantiate the predictive rule with TabPFN (Hollmann et al., 2022; 2025), a transformer trained on a broad range of synthetic tabular data that now achieves state-of-the-art results for tabular classification and regression tasks.
Our results show that TabMGP consistently provides reliable uncertainty quantification, often matching or surpassing both handcrafted martingale posteriors and classical Bayesian inference. These findings demonstrate that the MGP framework is considerably more robust than initially theorised: even when high-capacity predictive rules like TabPFN do not strictly satisfy formal sufficient conditions such as the martingale property, the resulting posteriors remain empirically stable, providing superior frequentist coverage and tighter credible sets than standard baselines. By establishing a new method for the modern Bayesian toolkit, this work opens new avenues for broadening the MGP framework to the era of foundation models.
2 Martingale posterior
The martingale posterior (Fong et al., 2023) is a recently proposed statistical inference method within the BPI framework (see Fortini and Petrone, 2025, for a review). We follow the notation of Fortini and Petrone (2025) in our discussion.
Notation. Let $(Z_{i})_{i\geq 1}$ be random variables taking values in a space $\mathcal{Z}$}. For any probability law $\mathbb{P}$} for the sequence $(Z_{i})_{i\geq 1}\in{\mathcal{Z}}^{\infty}$, define the predictive rule as the sequence of predictive distribution $P_{0}(\cdot)\coloneq{\mathbb{P}}(Z_{1}\in\cdot)$ and, for $i\geq 1$,
$P_{i}(\cdot)\coloneq{\mathbb{P}}(Z_{i+1}\in\cdot\mid Z_{1:i}).$
We write $P_{i}(\cdot\mid z_{1:i})$ for the realisation given $z_{1:i}$. We also write $\mathbb{P}$}-a.s. for “with $\mathbb{P}$}-probability one”. We use $p_{i}$ to denote the density or mass function of $P_{i}$.
The martingale posterior requires two ingredients:
  1. 1.
    A predictive rule $(P_{i})_{i\geq 0}$; and
  2. 2.
    A functional of interest $\theta:{\mathcal{P}}({\mathcal{Z}})\to\Theta\subseteq\mathbb{R}^{p}$, where ${\mathcal{P}}({\mathcal{Z}})$ denotes the set of all distributions $F$ on $\mathcal{Z}$}.
In this work, we focus on a risk minimiser that minimises a given loss function $\ell:{\mathcal{Z}}\times\Theta\to\mathbb{R}$:
$\theta(F)\coloneq\operatorname*{arg\,min}_{\vartheta\in\Theta}\int\ell(z,\vartheta)\,\mathrm{d}F(z)\,.$ (1)
For the martingale posterior distribution to be well-defined, the sequence of predictive distributions $(P_{i})_{i\geq 0}$ must converge $\mathbb{P}$}-a.s. to a random probability measure $F_{\infty}$, which represents the limiting empirical distribution of $(Z_{i})_{i\geq 1}$. Here, $F_{\infty}$ is analogous to the statistical model in classical Bayes, and the probability law of $F_{\infty}$ and $F_{\infty}\mid z_{1:n}$ are the prior and posterior respectively. This posterior law of $F_{\infty}\mid z_{1:n}$ then induces a posterior distribution on functionals of $F_{\infty}$ which is termed the martingale posterior.
Definition. Let $z_{1:n}\in{\mathcal{Z}}^{n}$ denote the observed data and $\Pi(F_{\infty}\mid z_{1:n})$ the posterior law of $F_{\infty}$ given $z_{1:n}$. The martingale posterior distribution is defined as the posterior law of $\theta(F_{\infty})$:
$\Pi_{\infty}(\theta(F_{\infty})\in A\mid z_{1:n})=\int{\bm{1}}(\theta(F_{\infty})\in A)\,\mathrm{d}\Pi(F_{\infty}\mid z_{1:n})\,$, (2)
for all measurable sets $A\subseteq\Theta$. The classical Bayesian posterior can be recovered from (2) as a special case by using a Bayesian PPD as the predictive rule and the corresponding negative log-likelihood as the loss function (Appendix B).
Sufficient conditions A sufficient, but not necessary, condition for the existence of $F_{\infty}$ and hence for the martingale posterior distribution to be well-defined, is that the sequence $(P_{i})_{i\geq 0}$ is a martingale, or equivalently, that the sequence $(Z_{i})_{i\geq 1}$ is conditionally identically distributed (c.i.d.) (Berti et al., 2004). The martingale property stipulates that $(P_{i})_{i\geq 0}$ satisfies, for every $i\geq 0$ and every measurable set $A$:
$\mathbb{E}[P_{i+1}(A)\mid Z_{1:i}]=P_{i}(A).$ (3)
Relaxations of this sufficient condition remain an active area of research (see, e.g., Battiston and Cappello, 2025). As noted in the introduction, the term “martingale posterior” is somewhat of a misnomer. Any predictive rule with a well-defined $F_{\infty}$ limit induces a posterior distribution as in (2). In this paper, we use the term in a broad sense: even predictive rules that violate the martingale property may still induce valid “martingale” posteriors.
Computation. The martingale posterior (2) generally does not admit a closed-form expression and is approximated by the finite martingale posterior. This is the posterior law of $\theta(F_{N})$, where $F_{N}$ is the empirical distribution of $Z_{1:N}$. A draw from $\theta(F_{N})$ is obtained by forward sampling a simulation rollout $Z_{n+1:N}$ from the predictive rule:
$Z_{i+1}\sim P_{i},\quad i=n,\dots,N-1,$
then forming $F_{N}$ and computing $\theta(F_{N})$ (Fong et al., 2023). Repeating this produces $L$ samples $\{\theta^{(l)}\}_{l=1}^{L}$. These rollouts are independent and thus embarrassingly parallel.
3 TabMGP
Given the strong predictive performance of modern foundation models (Bommasani et al., 2022), it is natural to ask whether such models can serve as a predictive rule in the martingale posterior framework. We describe our proposed methodology which employs TabPFN (Hollmann et al., 2022; 2025) as the predictive rule.
Given that TabPFN is designed for supervised prediction, we specialise the martingale posterior framework to the supervised setting, where each observation is a feature–response pair $z=(x,y)$. In this setting, a predictive rule generates future pairs $Z_{n+1},\ldots,Z_{N}$. Fong et al. (2023) describe this as the “joint method,” which requires modelling both the distribution of $X$ and the conditional distribution $Y\mid X$. Because modelling $X$ can be challenging even in moderate dimensions, we follow Fong et al. (2023) and use the Bayesian bootstrap (see Section 4) to forward sample features, drawing $X_{i+1}$ at random from the empirical distribution of $x_{1:i}$. In our work, we adopt this strategy for $X$, while TabPFN supplies the conditional $Y\mid X$.
3.1 Proposed methodology
We propose instantiating the martingale posterior with the predictive rule induced by TabPFN through in-context learning. TabPFN is a prior-data fitted network (Müller et al., 2022), explicitly trained to approximate a Bayesian PPD by sampling training data from exchangeable sequences defined by likelihood–prior pairs; see Appendix A for a review. We term the resulting martingale posterior TabMGP.
Several distinctive features of TabPFN facilitate its integration as a predictive rule within the martingale posterior framework. Through in-context learning, TabPFN consumes the context alongside a new query to return a one-step-ahead predictive distribution without the need for retraining or fine-tuning. This allows the model to adapt to novel datasets instantaneously, handling classification via deterministic probability vectors and regression through a multi-class classification approach that bins the response variable.
The architecture further benefits from native row-permutation invariance, ensuring that predictions depend solely on the multiset rather than the sequence order. This built-in property provides a significant computational advantage over many existing martingale posteriors, such as copula-based constructions (Fong et al., 2023), which must manually enforce invariance by averaging over permutations. Finally, as a prior-data fitted network, TabPFN is designed to be approximately Bayesian, having been trained to approximate Bayesian PPD. In the martingale posterior framework, the Bayesian PPD serves as the ideal predictive rule; when they coincide, the martingale posterior recovers the classical Bayesian posterior. By training on synthetic data from a vast array of likelihood–prior pairs, TabPFN is encouraged to approximate this ideal behaviour across a highly diverse range of tasks.
Why not other transformers? Large language models and other pretrained transformers often exhibit in-context learning: given a context $z_{1:i}$ and query $x_{i+1}$, they can produce a prediction for $y_{i+1}$. In this sense, they share the same basic mechanism as TabPFN. However, they lack the two further properties that make TabPFN especially well-suited as a predictive rule for martingale posteriors.
First, they are not row-permutation invariant. Since such models rely on positional embeddings, their predictions change with the ordering of training examples, unless explicit averaging is performed across permutations. Second, they are not trained to approximate Bayesian PPD: their training objectives are tailored to next-token prediction in text rather than calibrated probabilistic prediction for supervised tabular tasks. These differences mean that, although generic transformers can perform in-context learning, their predictive distributions are not naturally aligned with the martingale posterior framework. By contrast, TabPFN combines in-context learning with built-in permutation invariance and Bayesian alignment, making it a more principled choice.
Require : Observed data $z_{1:n}$, loss function $\ell(z,\theta)$, forward sampling depth NN, predictive rule $\mathrm{TabPFN}(\cdot\mid x_{i+1},z_{1:n})$, number of posterior samples $L$
for $l=1$ to $L$ do
 for $i=n$ to N1N-1 do
      Sample $x_{i+1}^{(l)}\leftarrow\operatorname{Empirical}(x_{1:i})$;
      Sample $y_{i+1}^{(l)}\leftarrow\mathrm{TabPFN}(\cdot\mid x_{i+1},z_{1:n})$;
      Form $z^{(l)}_{i+1}\leftarrow(x^{(l)}_{i+1},y^{(l)}_{i+1})$;
   end for
  Compute $\theta^{(l)}\leftarrow\operatorname*{arg\,min}_{\theta}\sum_{i=1}^{N}\ell(z_{i}^{(l)},\theta)$;
end for
Return : $\{\theta^{(1)},\dots,\theta^{(L)}\}$ as samples from the martingale posterior.
Algorithm 1 TabMGP
3.2 Validity of TabMGP.
A central question for any MGP construction is whether the chosen predictive rule leads to a well-defined limiting law $F_{\infty}$. While the framework traditionally points to conditions like the martingale property (Fong et al., 2023) or the almost conditionally identically distributed (a.c.i.d.) condition (Battiston and Cappello, 2025) as sufficient guarantees, verifying these for high-capacity models like TabPFN is currently impossible with existing tools in deep learning theory. Preliminary visual inspections of the predictive rule associated to TabPFN suggest there may be subtle deviations from these sufficient conditions (see Appendix C). However, we argue that such properties are sufficient rather than necessary for delivering useful epistemic uncertainty.
In the absence of formal theoretical guarantees, we rely on empirical diagnostics to verify the behaviour of TabMGP:
  • Path Stability (Scaled $L_{1}$ Convergence): The stability of the functional $\theta(F_{N})$ as the forward sampling depth $N$ increases is the primary diagnostic for the existence of a stable limit. We monitor this convergence by tracking the expected $L_{1}$-norm between the initial empirical risk minimiser $\theta(F_{n})$ and the terminal estimate $\theta(F_{N})$, scaled by the parameter dimension $p$: $\mathbb{E}_{F_{N}}[\frac{1}{p}\lVert\theta(F_{n})-\theta(F_{N})\rVert_{1}]$. If the transformer is internalising a consistent law, these trajectories should stabilise at a non-zero constant as $N$ grows. Systematic drift or divergent $L_{1}$ paths would indicate that there is no limiting measure.
  • Frequentist Coverage Assessment: We use frequentist coverage as an empirical check on the relationship between the transformer’s predictive rule and the population $F^{\star}$. A $(1-\alpha)$ credible set should contain the population risk minimiser $\theta(F^{\star})$ at a rate of at least $1-\alpha$.
  • Posterior Contraction: A valid epistemic uncertainty measure must represent uncertainty that vanishes with increasing data. One can verify this behaviour by observing the width of the posterior distribution of $\theta(F_{N})$ as the observed sample size $n$ increases. A sensible predictive rule should result in a distribution that narrows toward the population risk minimiser $\theta(F^{\star})$ as evidence accumulates, reflecting a reduction in epistemic uncertainty. See Figure 2 for an example.
Figure 2: Concentration of TabMGP. The black vertical line indicates the population risk minimiser $\theta(F^{\star})=(0,{\beta^{\star}}^{\top})$. See Appendix C for experimental details.
By focusing on these observed behaviours, practitioners can validate the outputs of TabMGP based on its empirical consistency, bypassing the currently insurmountable challenge of proving the transformer’s limiting properties.
4 Related works
The martingale posterior framework admits many possible predictive rules, and a growing literature has explored different constructions. We review these constructions below, since they form the main competing methods for TabMGP in Section 5. In parallel, a separate line of work has asked whether the in-context learning behaviour of transformers satisfies martingale properties, which is especially relevant here since TabPFN is a transformer foundation model whose in-context learning defines our predictive rule.
Recent analyses (Falck et al., 2024; Ye and Namkoong, 2024; Nagler and Rügamer, 2025) show that the predictive rules induced by foundation models, such as TabPFN, operate outside the traditional martingale property. Building on this line of work, we take the further step of assessing the more flexible almost conditionally identically distributed (a.c.i.d.) condition (Battiston and Cappello, 2025) and observe that TabPFN departs from this requirement as well. Nevertheless, we find that the resulting TabMGP exhibits consistent path stability and near-nominal coverage in practice. This suggests that the usual sufficient conditions are stricter than necessary and that validity can be achieved under a more relaxed set of assumptions than previously theorised.
Interestingly, Nagler and Rügamer (2025) adopt an approach closely aligned with ours, employing TabPFN as the predictive rule in the martingale posterior framework. Their finding that TabPFN is not a martingale leads them to abandon it as a predictive rule after the initial step, instead using its output to initialise a copula-based update that preserves martingale validity. By contrast, we retain TabPFN as the predictive rule throughout and interpret its departure from both the martingale and a.c.i.d. conditions as evidence that validity can be achieved under weaker assumptions.
A concurrent effort in Fortini et al. (2026) focuses on posterior inference for $F_{\infty}\mid z_{1:n}$ using a predictive central limit theorem (CLT). However, their framework does not incorporate the generalised Bayes flavour of inference for a parameter $\theta$ defined by a user-specified loss function. The predictive CLT has distinct advantages in supervised settings, and extending the predictive CLT to functionals of $\theta(F_{\infty})$ would make for interesting future work.
We now give a brief summary of the main predictive rules developed in the martingale posterior literature.
Bayesian boostrap The Bayesian bootstrap (Rubin, 1981) is a predictive rule where future observations $Z_{i+1}$ are uniformly drawn from the observations up to $i$:
$Z_{i+1}\sim F_{i},\quad i\geq n,$
with $F_{i}$ the empirical distribution of $z_{1:i}$. As this rule is discrete, its use is recommended only when the functional of interest arises from a parametric likelihood loss, for which the smoothness of $F_{\infty}$ is irrelevant (Fong et al., 2023; Dellaporta et al., 2022).
Newton’s algorithm. Rooted in Bayesian nonparametrics, Newton’s algorithm (Newton et al., 1998) was initially developed for unsupervised sequential learning in mixture models. It was subsequently re-interpreted as a predictive rule that generates a c.i.d. sequence of observations (Fortini and Petrone, 2020). This approach, however, is not suitable for high-dimensional $\theta$, as it requires numerical integration of an intractable normalising constant.
Copula-based updates. For applications requiring smoothness of $F_{\infty}$, the mainstream implementation of the martingale posterior is dominated by copula-based updates. Introduced by Hahn et al. (2018) and adopted as the default predictive rule in Fong et al. (2023), these updates smooth $F_{N}$ by attaching a bivariate copula to each new observation while maintaining the martingale property on the predictive rule exactly. Subsequent work has extended this framework with autoregressive Gaussian-process likelihoods for higher-dimensional features (Ghalebikesabi et al., 2023), direct quantile updates (Fong and Yiu, 2024b), log-concave shape constraints (Cui and Walker, 2024), and vine copulas for very high dimensional features (Huk et al., 2024). All these variants preserve the exact martingale guarantee of the copula framework, but each introduces at least one smoothing or dependence hyperparameter, whose value controls the spread of the resulting posterior. These hyperparameters are generally difficult to tune and must be retuned for each new dataset.
Parametric plug-in predictives. For parametric inference, Walker (2022) and Holmes and Walker (2023) proposed a predictive rule of the form $Z_{i+1}\sim p(\cdot\mid\widehat{\theta}_{i})$, where $p(\cdot\mid\theta)$ is an assumed parametric model and $\widehat{\theta}_{i}$ is a point estimate (e.g., maximum likelihood estimate) computed from all previous observations $z_{1:i}$. The asymptotic properties of this predictive rule are studied in Fong and Yiu (2024a). However, it requires that the assumed parametric model provides a reasonable approximation to the data distribution in order to yield sensible uncertainty quantification.
Permutation-equivariant neural network. Martingale posteriors have also been explored in the context of neural processes (Lee et al., 2023), where a permutation-equivariant neural network is used to construct a martingale posterior over latent variables. While this approach demonstrates the potential of using neural networks within a martingale posterior framework, the focus is still on prediction instead of inference, and it remains unclear how to generalise to broader statistical inference tasks.
5 Experiments
We compare TabMGP with other martingale posterior constructions using synthetic and real-world data. Our experiments focus on small- to moderate-$n$ settings where the prior specification of the classical Bayesian posterior is important, making the MGP particularly well suited.
To evaluate a given posterior $\Pi(\cdot\mid z_{1:n})$, we compute the coverage and ‘size’ of its joint $(1-\alpha)$ credible set $C_{1-\alpha}(z_{1:n})\subset\Theta$:
$\Pr_{\theta\sim\Pi(\cdot\mid z_{1:n})}\bigl(\theta\in C_{1-\alpha}(z_{1:n})\bigr)=1-\alpha.$
Although the set can be constructed in different ways, an ellipsoidal approximation is often used for computational efficiency; see Appendix D for a discussion.
The coverage of $C_{1-\alpha}(z_{1:n})$ is defined as
$\Pr_{z_{1:n}\sim F^{\star}}\bigl(\theta(F^{\star})\in C_{1-\alpha}(z_{1:n})\bigr),$ (4)
namely, the proportion which the credible set contains the population risk minimiser, $\theta(F^{\star})=\arg\min_{\theta}\int\ell(z,\theta)\,\mathrm{d}F^{\star}(z)$, under repeated draws of $z_{1:n}$ from the true data-generating distribution $F^{\star}$. Although frequentist in nature, this metric is commonly used for comparing Bayesian posteriors.
Posterior constructions. We compare TabMGP against the martingale posteriors suggested in Fong et al. (2023): Bayesian bootstrap (BB) and the bivariate copula update (Copula). As a baseline, we also include the classical Bayesian posterior (Bayes) using asymptotic approximation of the posterior as the prior. We also compare TabMGP against this asymptotic approximation of the posterior (Asymptotic). The model choice is discussed later in this section when introducing the loss functions for this experiment. The exact computation details of the posteriors are provided in Appendix E.1.
Loss functions. We use negative log-likelihood as the loss in (1), and the corresponding model in the Bayes posterior. For continuous response (linear regression), we use $\ell(x,y,\theta)=(y-[1\ x^{\top}]\theta)^{2}$ which is the negative log-likelihood of ${\mathcal{N}}([1\ x^{\top}]\theta,1)$. For $K$-class categorical response (logistic regression), we use $\ell(x,y,\theta)=-\log\Pr(y=k)$, where $\Pr(y=k)$ is the $k$-th entry of the softmax applied to the logits $([1\ x^{\top}]\theta_{1},\ldots,[1\ x^{\top}]\theta_{K})$, and $\theta_{1},\ldots,\theta_{K}$ are the class-specific coefficients. We also set appropriate constraints on the coefficients to ensure identifiability in the loss function; see Appendix E.2 for details.
Data-generating setups. We describe the true data-generating distribution $F^{\star}$ of this experiment. For each $F^{\star}$, we repeatedly sample 100 repetitions $\{z_{1:n}^{(r)}\}_{r=1}^{100}$ from $F^{\star}$. We refer to $F^{\star}$ together with the procedure used to generate $z_{1:n}$ as a setup. The $n$ is chosen to be as small as possible without rendering the design matrix ill-conditioned. In total, we have 30 setups (11 synthetic, 19 real-data).
Let $\beta^{\star}$ be a set of regression coefficients and $x_{i}$ i.i.d. samples from a hypercube $1,1]^{10}$. For the synthetic setups, we generate $y\mid x$ with:
  1. 1.
    $y_{i}=x^{\top}_{i}\beta^{\star}+\epsilon_{i}$ where $\epsilon_{i}\overset{iid}{\sim}{\mathcal{N}}(0,1)$;
  2. 2.
    $y_{i}=x^{\top}_{i}\beta^{\star}+\epsilon_{i}$ where $\epsilon_{i}$ are i.i.d. and follows a Student’s t-distribution: $t_{5}$, $t_{4}$ or $t_{3}$;
  3. 3.
    $y_{i}=x^{\top}_{i}\beta^{\star}+\epsilon_{i}$ where $\epsilon_{i}$ is heteroscedastic and depends on $x_{i}$. We have three different strength of heteroscedasticity: $s_{1}$, $s_{2}$, $s_{3}$, from weak to strong;
  4. 4.
    $y_{i}\mid x_{i}\overset{\mathrm{iid}}{\sim}\mathrm{Bernoulli}(L(x^{\top}_{i}\beta^{\star}))$, with a logistic link function $L$;
  5. 5.
    $y_{i}\mid x_{i}\overset{\mathrm{iid}}{\sim}\mathrm{Bernoulli}(L(x^{\top}_{i}\beta^{\star}))$, where $L$ is the distribution function of a Gaussian mixture model (GMM). We consider three choices of location parameters: GMM(0)(0), GMM(1)(-1), and GMM(2)(-2), where the argument denotes the mean of one mixture component.
For the 19 real-data setups, we use datasets from the OpenML and UCI repositories (9 continuous response, 8 binary response, and 2 multiclass categorical response). As $F^{\star}$ is unknown in the real-data setups, we take the empirical distribution of the entire dataset as the truth $\widetilde{F}^{\star}$. The details of all data-generating setups are given in Appendix E.3.
Metrics for comparing posteriors. For each posterior, we construct joint credible sets $\widehat{C}_{1-\alpha}(z_{1:n})$ and compare their coverage (4) and ‘size’. Our approach of constructing $\widehat{C}_{1-\alpha}(z_{1:n})$ is described in Appendix E.4. We consider a posterior more reliable and meaningful when its credible set is small while achieving coverage at least $1-\alpha$.
In the synthetic setups, the coverage is given by $\frac{1}{100}\sum^{100}_{r=1}{\bm{1}}\bigl[\theta(F^{\star})\in\widehat{C}_{1-\alpha}(z^{(r)}_{1:n})\bigr$], where $\{z_{1:n}^{(r)}\}_{r=1}^{100}$ are 100 replicates of $z_{1:n}$ drawn from $F^{\star}$. In the real-data setups, as $F^{\star}$ is unknown, we take the empirical distribution of the entire dataset $\mathcal{D}$} as the truth $\widetilde{F}^{\star}$, and compute the coverage on $\theta(\widetilde{F}^{\star})=\arg\min_{\theta}\sum_{z\in{\mathcal{D}}}\ell(z,\theta)$. That is, the coverage is $\frac{1}{100}\sum^{100}_{r=1}{\bm{1}}\bigl[\theta(\widetilde{F}^{\star})\in\widehat{C}_{1-\alpha}(z^{(r)}_{1:n})\bigr$], where $z^{(r)}_{1:n}$ denotes the $r$-th random draws (without replacement) from $\mathcal{D}$}; see Appendix E.3 for details. To quantify the ‘size’ of a credible set, we use the trace of the posterior covariance matrix as a proxy.
5.1 Results
Convergence of $\theta(F_{N})$. To conduct statistical inference, we require $\theta(F_{N})$ to converge $\mathbb{P}$}-a.s. as $N\to\infty$. This convergence is commonly assessed using trace plots that track each component of $\theta(F_{N})$ as $N$ increases, but this approach quickly becomes impractical in high dimensions. Instead, we track the expected $L_{1}$-norm of $\theta(F_{N})$ relative to the empirical risk minimiser $\theta(F_{n})$ as $N$ increases; see Appendix F for the precise definition. For convergent $\theta(F_{N})$, we desire to see the expected $L_{1}$-norm stabilise at a non-zero constant as $N$ increases.
We present the trajectories in Figure 3, where each trajectory corresponds to one of the 30 setups. For completeness, we also report the $L_{1}$-norm between each realisation $F_{N}$ and $F_{n}$ in Figure 6. In general, we observe that $N=500$ is sufficient for convergence in most setups, while some setups are near convergence and are expected to converge as the number of forward sampling steps increases.
Figure 3: Expected $L_{1}$-norm between $\theta(F_{N})$ from TabMGP and $\theta(F_{n})$ as $N$ increases. Each of the 30 trajectories corresponds to a realisation of $z_{1:n}$ from one setup.
Shape of the posteriors. We present the marginals of the posteriors for a realisation of $z_{1:n}$ in Figure 4 (intercept of the real-data ‘concrete’ setup), with the remaining dimension and setups provided in Appendix G. In general, the posterior means are similar across methods, but TabMGP often exhibits skewness and multimodal structure compared with the other posteriors, which are typically Gaussian-like. This suggests that TabMGP is more expressive.
Figure 4: Posterior densities of the intercept in the ‘concrete’ setup. For each posterior density, the 95% marginal credible interval is shown as a horizontal bar, and the posterior mean is marked. The solid and dashed black vertical lines correspond to $\theta(F^{\star})$ and $\theta(F_{n})$ respectively.
Joint credible sets of linear regression. The results for synthetic data (top and middle) and real-data (bottom) setups are shown in Table 1. In the real-data, the rows are ordered by the $n/p$ ratio in descending order, with $p\coloneq\dim(\theta)$.
TabMGP is competitive in both coverage and credible set size, performing particularly well on real data. Notably, it can attain nearly 100% coverage in many setups while maintaining small credible sets. Similarly, Copula provides good coverage with small credible sets, although its performance is less consistent across real-data setups (e.g., severe undercoverage in kin8nm and overcoverage in quake). This is expected, as bespoke Gaussian-copula-based predictive rules have a structural advantage over TabMGP in Gaussian (or near-Gaussian) setups. Bayes and Asymptotic produce excessively large credible sets in all synthetic setups, as both rely on asymptotic approximations that are less accurate in low-$n$ settings. They are also more susceptible to model misspecification and tend to underperform relative to TabMGP. In contrast, BB tends to produce overly small sets with poor coverage in the synthetic setups (top and middle) and in low $n/p$ settings (last few rows of Table 1). This reflects limited variability in forward samples when bootstrapping from a small $z_{1:n}$ (e.g., $n=20$ in the synthetic data setups).
Table 1: Experiments on linear regression: i.i.d. errors (top), dependent errors (middle), and real data (bottom). We compare TabMGP (ours), BB, Copula, Bayes and Asymptotic. The rows corresponding to real data are sorted by $n/p$ ratio, i.e, bottom rows are harder. Reported values are coverage (Rate) and the median of trace of the posterior covariance (Size), computed over 100 repetitions. Target coverage is 0.95.
TabMGP $B$ Copula Bayes Asymptotic
Setup Rate Size Rate Size Rate Size Rate Size Rate Size
${\mathcal{N}}(0,1)$ 1.00 0.45 0.55 0.09 0.99 0.35 1.00 0.65 1.00 1.31
$t_{5}$ 1.00 0.49 0.65 0.11 0.99 0.37 1.00 0.65 1.00 1.31
$t_{4}$ 0.99 0.51 0.64 0.12 0.98 0.37 0.99 0.65 1.00 1.31
$t_{3}$ 1.00 0.48 0.66 0.14 0.97 0.35 0.98 0.65 0.98 1.31
$s_{1}$ 1.00 0.37 0.50 0.03 1.00 0.36 1.00 0.65 1.00 1.31
$s_{2}$ 1.00 0.36 0.46 0.02 1.00 0.36 1.00 0.65 1.00 1.31
$s_{3}$ 1.00 0.33 0.53 0.02 1.00 0.37 1.00 0.65 1.00 1.31
concrete 0.91 0.06 0.80 0.05 1.00 0.12 0.87 0.05 1.00 0.10
quake 1.00 0.03 0.95 0.07 1.00 0.18 0.91 0.04 0.98 0.09
airfoil 0.96 0.08 0.93 0.05 0.97 0.11 0.96 0.06 1.00 0.12
energy 1.00 0.04 0.80 0.01 1.00 0.06 1.00 0.07 1.00 0.14
fish 0.90 0.10 0.79 0.07 0.95 0.10 0.91 0.08 1.00 0.16
kin8nm 0.95 0.13 0.77 0.09 0.63 0.08 0.84 0.11 0.99 0.22
auction 0.99 0.66 0.86 0.35 0.97 1.44 1.00 0.81 1.00 1.65
grid 0.97 0.13 0.72 0.08 0.88 0.10 0.97 0.16 1.00 0.31
abalone 0.97 0.95 0.66 0.69 0.94 0.98 0.84 0.81 0.98 1.64
Joint credible sets of logistic regression. The results for synthetic (top) and real-data (middle and bottom) setups are in Table 2. Copula was not applied to multinomial logistic regression (bottom) as the implementation of Fong et al. (2023) is not directly applicable. In the real-data setups, the rows are ordered by decreasing $n/p$ within each section.
In the synthetic setups, the Asymptotic credible sets achieve the best coverage and size, as expected, since the asymptotic approximation is reasonably accurate at the larger sample size ($n=100$). TabMGP, however, shows the most consistent performance in the real-data setups, often achieving good coverage with the smallest credible sets. TabMGP performs particularly well relative to competing methods in the ‘harder’ setups, i.e., the bottom rows with low $n/p$ ratios. Asymptotic also provides reasonable coverage in the real-data setups, although its credible sets are typically much wider than those of TabMGP. This is likely due to model misspecification, which results in miscalibrated credible sets. BB performs reasonably well in binary logistic regression (top and middle), but its credible sets are larger than those of TabMGP. However, BB severely under-covers in multinomial logistic regression (bottom), where $p$ is much larger than in the other logistic regression setups. Copula performs poorly in logistic regression, producing trivial credible sets. This suggests that Copula is sensitive to the choice of bandwidth, as the same value yields good performance in the linear regression experiment.
Table 2: Experiments on logistic regression: synthetic data with various link functions (top), real data with binary logistic regression (middle), and real data with multinomial logistic regression (bottom). The rows corresponding to real data are sorted by $n/p$ ratio, i.e, bottom rows are harder. We compare TabMGP (ours), BB, Copula, Bayes and Asymptotic. Reported values are coverage (Rate) and the median of trace of the posterior covariance (Size), computed over 100 repetitions. Target coverage is 0.95.
TabMGP $B$ Copula Bayes Asymptotic
Setup Rate Size Rate Size Rate Size Rate Size Rate Size
Logistic 0.84 1.30 0.83 1.96 1.00 2.72 0.53 0.78 0.96 1.53
GMM(0) 0.91 2.44 0.91 4.14 1.00 4.01 0.60 1.34 0.98 2.63
GMM(-1) 0.83 1.27 0.86 1.95 1.00 2.64 0.58 0.79 0.99 1.54
GMM(-2) 0.78 0.88 0.78 1.36 1.00 2.08 0.52 0.57 1.00 1.13
rice 1.00 3.74 0.88 7.35 1.00 28.41 0.91 2.87 1.00 5.92
sepsis 0.95 0.69 1.00 2.42 1.00 10.00 0.98 1.18 1.00 2.42
banknote 0.99 1.16 1.00 1.48 1.00 1.53 0.99 0.64 1.00 1.28
mozilla 0.96 1.42 0.97 2.04 1.00 2.68 0.87 0.65 0.93 1.31
skin 1.00 1.39 1.00 1.84 1.00 1.80 1.00 0.57 1.00 1.16
blood 0.87 1.07 0.95 2.51 1.00 5.40 0.84 0.86 0.99 1.71
phoneme 1.00 1.32 0.95 1.92 1.00 5.62 0.73 0.83 1.00 1.66
telescope 0.99 5.96 0.99 10.66 1.00 50.69 0.67 3.50 0.96 7.10
yeast 0.97 11.10 0.68 25.13 0.33 10.82 0.88 19.22
wine 0.93 11.04 0.79 39.75 0.09 12.51 0.91 26.25
Marginal credible intervals. In addition to the results of joint credible sets presented above, we also report the results of marginal credible intervals in Appendix H. We report both the coverage and the median width of the credible intervals.
Computation time. The main computational cost of martingale posteriors is forward sampling, which is readily parallelisable. Per sample, BB and Asymptotic are the fastest (near-instantaneous); Copula takes 10–20s; and TabMGP takes about 70s (binary/categorical response) or 200s (continuous response). However, despite being the most computationally expensive method, TabMGP requires minimal manual intervention compared to the others. The NUTS algorithm (Bayes) runs in under 60s. BB, Copula, Bayes, and Asymptotic are implemented in compiled JAX, whereas TabMGP currently runs in uncompiled PyTorch. We expect this gap to narrow with further code optimisation. All experiments were conducted on an Nvidia L40s.
6 Conclusion
While bespoke Bayesian modelling remains the gold standard when a well-specified likelihood–prior pair is available, tabular foundation models like TabPFN are increasingly used in practice. In scientific workflows, prediction is usually a means to an end: one ultimately wants uncertainty about a scientific estimand $\theta$. However, TabPFN does not by itself return posterior uncertainty for $\theta$. In TabPFN, the carrier of epistemic uncertainty is the predictive rule itself and the associated random limiting measure $F_{\infty}$, so the natural inference target is the functional $\theta(F_{\infty})$. This motivates a framework that returns posterior uncertainty for $\theta(F_{\infty})$ using only queries to TabPFN’s predictive outputs.
For this purpose, we introduced TabMGP. The procedure requires neither likelihood–prior specification, as in classical Bayesian inference, nor bespoke predictive rules or hyperparameter tuning, as in existing martingale posterior frameworks, relying instead solely on the predictive rule induced by TabPFN. We evaluated TabMGP across a large collection of synthetic setups and real-world datasets, and compared extensively to existing martingale posterior constructions as well as standard Bayesian baselines. Across these experiments, TabMGP was consistently favourable in terms of frequentist behaviour, delivering stable posteriors with near-nominal coverage. Our empirical evaluation shows that TabMGP provides reliable uncertainty quantification.
Our findings also expose limitations in the current theoretical understanding of martingale posteriors. Existing analyses emphasise sufficient conditions such as the martingale or a.c.i.d. properties; in our experiments, finite-horizon diagnostics do not certify these sufficient conditions, yet the resulting posteriors maintain empirical stability and near-nominal coverage. This suggests that the present theory may be overly restrictive and does not yet capture the range of predictive rules that work well empirically. Developing weaker and more realistic conditions is an important direction for future research, particularly as martingale posteriors are likely to be combined with an increasing variety of black-box predictive models. More broadly, deep learning theory continues to lag far behind practice: for predictors like TabPFN, we should not expect to be able to theoretically verify that any proposed condition holds, so progress will also require a battery of theory-guided empirical diagnostics to test these conditions.
Acknowledgments
The authors would like to thank Abdelhamid Ezzerg, Joshua Bon, Christian Robert, Lorenzo Cappello, Jack Jewson, Rubén Loaiza-Maya and Anastasios Panagiotelis for helpful comments in the development of this work. KN was supported by the Australian Government Research Training Program and the Statistical Society of Australia PhD Top-up Scholarship. EF was supported by the Research Grants Council of Hong Kong through the General Research Fund (17307321) and the Early Career Scheme (27304424). DTF acknowledges funding from the Australian Research Council (DE200101070, DP200101414). JK was supported through the UK’s Engineering and Physical Sciences Research Council (EPSRC) via EP/W005859/1. SW was supported by the Australian Research Council (DE200101253). Computational resources were provided by the ARDC Nectar Research Cloud.
Appendix
Appendix A Review of TabPFN
A precursor to TabPFN appears in Müller et al. (2022), where the transformer is referred to as a prior-data-fitted network (PFN). The underlying premise is that the Bayesian posterior predictive distribution (PPD) is optimal in an average sense. We present this result in the unsupervised setting to keep the notation lightweight, drawing on a result from Aitchison (1975) [^1] The work of Müller et al. (2022) does not appear to be aware of this result..
Suppose we have an exchangeable distribution
$p(z_{1},\ldots,z_{n})=\int\prod_{i=1}^{n}p(z_{i}\mid\nu)\phi(\nu)\,\mathrm{d}\nu\,.$ (5)
Consider the Kullback-Leibler (KL) divergence between $p(z\mid\nu)$ and some predictive density $q(z\mid z_{1:n})$ where $z_{1:n}$ is drawn from (5), i.e., first draw $\nu$ from the prior $\phi(\nu)$ and then draw $z_{i}$ i.i.d. from $p(z\mid\nu)$. Then,
$D_{\mathrm{KL}}(p(z\mid\nu)\,\|\,q(z\mid z_{1:n}))$
is random because a) $\nu$ is random and b) the data $z_{1:n}$ is random. By integrating out these sources of randomness, we get an average KL divergence, a deterministc quantity
$\mathbb{E}_{\nu,z_{1:n}}D_{\mathrm{KL}}(p(z\mid\nu)\,\|\,q(z\mid z_{1:n}))=\int\phi(\nu)\,\mathrm{d}\nu\int p(z_{1:n}\mid\nu)\,\mathrm{d}z_{1:n}\int p(z\mid\nu)\log\frac{p(z\mid\nu)}{q(z\mid z_{1:n})}\,\mathrm{d}z\,.$ (6)
Aitchison (1975) then shows that the predictive density $q$ that minimises this average KL divergence in (6) is none other than the Bayesian PPD corresponding to the likelihood–prior pair in (5):
$p(z\mid z_{1:n})=\int p(z\mid\nu)\pi(\nu\mid z_{1:n})\,\mathrm{d}\nu$ (7)
where $\pi(\nu\mid z_{1:n})\propto\prod_{i=1}^{n}p(z_{i}\mid\nu)\phi(\nu)$, the standard Bayesian posterior construction.
This optimality result for the Bayesian PPD immediately suggests a meta-learning training objective. Since minimising the average KL divergence with respect to $q$ is indifferent to the entropy term of $p(z\mid\nu)$ in the KL divergence, we may rewrite the objective as
$\operatorname*{arg\,min}_{q(z\mid z_{1:n})}E_{\nu,z_{1:n}}\operatorname{KL}(p(z\mid\nu)\,\|\,q(z\mid z_{1:n}))=\operatorname*{arg\,min}_{q(z\mid z_{1:n})}E_{\nu,z_{1:n}}E_{p(z\mid\nu)}-\log q(z\mid z_{1:n})$
to which the empirical risk counterpart is
$\operatorname*{arg\,min}_{w}-\sum_{m=1}^{M}\sum_{k=1}^{K}\log q_{w}(z_{n+k}^{m}\mid z_{1:n}^{m}).$
if we use a model $q_{w}$. Here, $q_{w}$ can be any model that can accept a variable length set $z_{1:n}$ and returns a distribution over future $z$, and the samples $z_{1:n+K}$ are drawn from an exchangeable distribution $p(z_{1},\ldots,z_{n},z_{n+1},\ldots,z_{n+K})$ as in (5). Müller et al. (2022) choose to use a variant of transformers for $q_{w}$. This is most natural when $z$ is categorical. In the case of real-valued $z$, Müller et al. (2022) perform an adaptive binning operation that essentially converts the regression to a classification task.
Later developments led to the TabPFN (Hollmann et al., 2022; 2025) which trains on a vast amount of synthetic data generated from a very large array of likelihood–prior pairs (e.g., structural causal models and Bayesian neural network models), rather than a single pair as in the original PFN work. This gives TabPFN the qualifier foundational.
Appendix B Recovering classical Bayesian posterior from martingale posterior.
A sample from the classical Bayesian posterior $\pi(\theta\mid z_{1:n})\propto\pi(\theta)\prod_{i=1}^{n}p(z_{i}\mid\theta)$ can be equivalently obtained from the martingale posterior. This is achieved by forward sampling with Bayesian PPD:
$p(z\mid z_{1:i})=\int p(z\mid\theta)\,\pi(\theta\mid z_{1:i})\,\mathrm{d}\theta,$
then compute the posterior mean estimator $\mathbb{E}[\theta\mid z_{1:\infty}]$. This is a result from the Doob’s martingale convergence theorem (Doob, 1949). Alternatively, we can use the maximum likelihood estimator as our functional of interest $\theta(F)$, i.e., setting the loss in (1) as $\ell(z,\theta)=-\log p(z\mid\theta)$, and the distribution of $\theta(F_{\infty})$ will be $\pi(\theta\mid z_{1:n})$. This is due to the asymptotic equivalence of a posterior mean estimator and the maximum likelihood estimator.
Appendix C Validity of TabMGP as a martingale posterior.
As discussed in Section 2, a sufficient condition for the existence of $F_{\infty}$ and hence for the martingale posterior distribution to be well-defined, is that the sequence $(P_{i})_{i\geq 0}$ is a martingale, or equivalently, that the sequence $(Z_{i})_{i\geq 1}$ is conditionally identically distributed (c.i.d, Berti et al., 2004). The martingale property stipulates that $(P_{i})_{i\geq 0}$ satisfies, for every $i\geq 0$ and every measurable set $A$:
$\mathbb{E}(P_{i+1}(A)\mid Z_{1:i})=P_{i}(A).$ (8)
A relaxations of this sufficient condition is the almost conditionally identically distributed (a.c.i.d.) condition (Battiston and Cappello, 2025). Let $(\xi_{i})_{i\geq 1}$ be a sequence of non-negative random variables. The sequence $(Z_{i})_{i\geq 1}$ is a.c.i.d. if the corresponding predictive rule $(P_{i})_{i\geq 0}$ satisfies
$\lvert\mathbb{E}_{Z_{i+1}}(P_{i+1}(A)\mid Z_{1:i})-P_{i}(A)\rvert\leq\xi_{i},\quad\text{a.s.,}$ (9)
for all $i\geq 1$ and all measurable sets $A$. Battiston and Cappello (2025) show that $F_{\infty}$ exists if $\sum_{i=0}^{\infty}\xi_{i}<\infty$, $\mathbb{P}$}-a.s., and that the sequence $(Z_{i})_{i\geq 1}$ is $F_{\infty}$-asymptotically exchangeable. We will refer to these conditions collectively as the a.c.i.d. condition. Clearly, (9) generalises the stronger martingale condition (8), i.e., $\xi_{i}=0$ for all $i\geq 1$, and we will use (9) to verify the existence of $F_{\infty}$ for TabMGP instead of (8).
Verifying (9) empirically is generally challenging, as it must hold for all measurable sets $A$. Rather, we adopt an equivalent formulation in terms of total variation distance (Battiston and Cappello, 2025):
$\sum_{i=1}^{\infty}\operatorname{TV}\left(\mathbb{E}_{Z_{i+1}}(P_{i+1}(\cdot)\mid Z_{1:i}),P_{i}(\cdot)\right)\leq\xi_{i},\quad\text{a.s.,}$
for all $i\geq 0$. Since the predictive distributions of both TabPFN and the Bayesian bootstrap are probability mass functions $p_{i}$, the condition can be further simplified in terms of the $L_{1}$ distance:
$\sum_{i=1}^{\infty}\sum_{z\in{\mathcal{Z}}}\lvert\mathbb{E}_{Z_{i+1}}[p_{i+1}(z)\mid Z_{1:i}]-p_{i}(z)\rvert<\infty,\quad\text{a.s.}.$ (10)
There are several practical considerations when estimating (10). First, we start the outer summation from $i=n$, as TabPFN requires a small number of initial observations to produce reliable predictions. Second, in our supervised learning setting where $z=(x,y)$, the expectation
$\mathbb{E}_{Z_{i+1}}[p_{i+1}(z)\mid z_{1:i}]=\mathbb{E}_{Y_{i+1},X_{i+1}}[p_{i+1}(y\mid x_{i+2})\,p_{i+1}(x)\mid z_{1:i}]\,$,
requires a large number of Monte Carlo samples and is computationally demanding. To mitigate this, we fix all future covariates $x_{n+1:N+2}$ to a constant value $x^{\star}$ and consider the weaker condition:
$\sum_{i=n}^{N}\sum_{y\in{\mathcal{Y}}}\lvert\mathbb{E}_{Y_{i+1}}[p_{i+1}(y\mid x^{\star})\mid z_{1:i}]-p_{i}(y\mid x^{\star})\rvert$$\infty$, (11)
for a sufficiently large $N$. We then compute (11) repeatedly over many realisations of $Y_{n+1:N}$.
We evaluate this condition on an initial dataset of size $n=100$ generated as follows: $x_{i}$ are drawn independently and uniformly from the two-dimensional cube $1,1]^{2}$, and $y_{i}\overset{\mathrm{i.i.d.}}{\sim}\operatorname{Bernoulli}(\operatorname{logistic}(x_{i}^{\top}\beta^{\star}))$, for $i=1,\ldots,n$, with a fixed $\beta^{\star}\in\mathbb{R}^{2}$. We consider 10 different values of $x^{\star}$, randomly selected from the initial sequence $x_{1},\ldots,x_{n}$. Due to computational constraints, we evaluate (11) only on a sparse grid $i\in\{100,120,140,\ldots,1100\}$ and use the trapezoidal rule to approximate the partial sum.
Our empirical results (Figure 5) suggest that the predictive rule induced by TabPFN exhibits a lack of strict adherence to the a.c.i.d. condition as $N$ increases. However, at these scales ($N>1000$), any measurable drift falls below the numerical resolution of the model, making it impossible to distinguish genuine theoretical departures from the cumulative numerical instability inherent to long-horizon autoregressive inference. This empirical ambiguity, however, does not imply non-convergence to $F_{\infty}$ for TabMGP, as the a.c.i.d. condition is sufficient but not necessary. Moreover, the observed divergence may be attributable to the finite numerical precision of TabPFN. In particular, both $\mathbb{E}_{Y_{i+1}}[p_{i+1}(y\mid x^{\star})\mid z_{1:i}]$ and $p_{i}(y\mid x^{\star})$ are expected to be accurate only up to machine precision, so that their difference behaves as random numerical noise for sufficiently large $N$, and the partial sum is infeasible to compute in this regime.
Figure 5: Cumulative sum (over NN) of the $L_{1}$-distance between $\mathbb{E}_{y_{i+1}}[p_{i+1}(y\mid x^{\star})\mid z_{1:i}]$ and $p_{i}(y\mid x^{\star})$ across ten choices of $x^{\star}$. Ideally, this cumulative sum converges to satisfy the a.c.i.d. condition.
As an alternative, we track the sample path of $\theta(F_{N})$ as $N$ increases, as in Fong et al. (2023). After all, we need $\theta(F_{N})$ to converge $\mathbb{P}$}-a.s. to conduct meaningful statistical inference. We refer to these diagnostic plots as the “trace plot” and inspect them in Appendix F. We observe that the sample path of $\theta(F_{N})$ does converge in most of our experiment, thus suggesting that the a.c.i.d. condition is too restrictive for our purpose of establishing the existence of $\theta(F_{\infty})$.
For completeness, we further demonstrate that TabMGP can concentrate close to $\theta(F^{\star})$ as the sample size $n$ grows. To illustrate this, we use a regression dataset, since the sum of the sample size needed for meaningful concentration and the forward sampling steps required for convergence in classification exceeds TabPFN’s 10,000-context length limit. The regression dataset is defined as $y_{i}\overset{i.i.d.}{\sim}{\mathcal{N}}(x^{\top}_{i}\beta^{\star},0.1^{2}),\,i=1,\dots,n$, for some fixed $\beta^{\star}\in\mathbb{R}^{2}$, with the loss function $\ell(x,y,\theta)=(y-[1\ x^{\top}]\theta)^{2}$. Ideally, TabMGP should concentrate on $\theta(F^{\star})=[0,{\beta^{\star}}^{\top}]$ as $n$ increases. Figure 2 confirms that TabMGP indeed concentrates around $\theta(F^{\star})$.
Appendix D Highest-posterior density credible sets
The highest-posterior-density (HPD) credible set is a common construction of a Bayesian credible set. The HPD set at level $1-\alpha$ is defined as
$C_{1-\alpha}^{\mathrm{HPD}}=\left\{\theta\in{\mathbb{R}}^{p}:\pi(\theta\mid z_{1:n})\geq t_{1-\alpha}\right\},$
where $t_{1-\alpha}$ is chosen such that $\int_{C_{1-\alpha}^{\mathrm{HPD}}}\pi(\theta\mid z_{1:n})\,\mathrm{d}\theta=1-\alpha$. By construction, $C_{1-\alpha}^{\mathrm{HPD}}$ has minimum volume among all sets of posterior probability $1-\alpha$, making it the “gold standard” for summarising joint uncertainty (Hyndman, 1996).
In many models — particularly linear and logistic regressions with well-behaved posteriors — the posterior is approximately multivariate normal:
[…]
In these cases, the HPD set reduces to the $\chi^{2}$-ellipsoid
$C_{1-\alpha}^{\mathrm{HPD}}\approx\Bigl\{\theta:(\theta-\hat{\mu})^{\top}\hat{\Sigma}^{-1}(\theta-\hat{\mu})\leq\chi^{2}_{p,1-\alpha}\Bigr\},$ (12)
treating the Mahalanobis ellipsoid as the HPD set under the normal approximation.
Appendix E Experimental settings and hyperparameters
E.1 Posterior constructions
Asymptotic approximation of the Bayesian posterior. Let $\widehat{\theta}_{n}=\operatorname*{arg\,min}_{\theta}\sum^{n}_{i=1}\ell(z_{i},\theta)$ and $\widehat{H}_{n}=\nabla^{2}_{\theta}\sum^{n}_{i=1}\ell(z_{i},\theta)|_{\theta=\widehat{\theta}_{n}}$. We use ${\mathcal{N}}(\widehat{\theta}_{n},\widehat{H}_{n}^{-1})$ as the asymptotic approximation of the posterior for both constructing the credible region and using it as the prior for the classical Bayes baseline.
Settings of Bayes. We approximate the posterior using a No-U-turn sampler (NUTS, Hoffman and Gelman, 2014) with 1000 warm-up steps, followed by 2000 sampling steps across 4 chains with different initialisations, for a total of 8000 samples.
Settings for all martingale posteriors. We run 2000 forward-sampling steps for both BB and Copula. For TabMGP, we find that $N=n+500$ is sufficient for $\theta(F_{N})$ to converge (Figure 3 and Appendix F) and use only 500 forward-sampling steps. We approximate all martingale posterior constructions with 100 samples.
Settings of TabMGP. The logits of the predictive rules are obtained from TabPFNClassifier.predict_proba for logistic regression and TabPFNRegressor.predict(output_type = "full") for linear regression. By default, these logits are scale by a temperature of 0.9. For our experiments, we use the untempered logits as our predictive rules, i.e., setting the temperature to 1. By construction, TabPFN is not permutation invariant to the order of the dimensions of labels $y$ and features $x$. The default implementation of TabPFN uses an ensemble of predictions from randomly permuting the dimensions of labels and features. The resulting probability mass functions are then averaged to get the final predictive mass function. We use the default number of ensemble in version 2.0.6, which is 4 for classifiers and 8 for regressors.
We use the following two checkpoints:
These are the default checkpoints at the time of running our experiments, and the transformer was pre-trained exclusively on synthetic, exchangable sequence of data. Subsequent iterations of the TabPFN checkpoints, however, have been fine-tuned on real-world datasets, and these checkpoints are not used in our experiments.
Settings of Copula. We use the copula construction in Fong et al. (2023, §4.4.2 and §4.4.3) with the bandwidth fixed to 0.8, since for our experiments the marginal-likelihood-based tuning procedure of Fong et al. (2023) produces values that are too small and delivers poor results. Instead, we estimate
$\theta(P_{N})\approx\operatorname*{arg\,min}_{\theta}\sum_{i}\ell(\widetilde{x}_{i},\widetilde{y}_{i},\theta),$
where $(\widetilde{x}_{i},\widetilde{y}_{i})_{i\geq 0}$ are generated by iterating $\widetilde{x}_{i}$ over $x_{1:n}$ and, for each $\widetilde{x}_{i}$, drawing $\widetilde{y}_{i}$ from $P_{N}(\cdot\mid\widetilde{x}_{i})$. We repeat this procedure five times, yielding $5n$ pairs of $(\widetilde{x}_{i},\widetilde{y}_{i})$.
E.2 Specification of the loss function
We set the following coefficients to 0 to ensure identifiablity in the loss function:
  1. 1.
    The coefficients for the first level of each categorical feature;
  2. 2.
    The coefficients corresponding to all (almost) collinear features;
  3. 3.
    In logistic regression, the coefficients corresponding to the first category, $\theta_{1}$.
We identify the collinear features as those with the largest loading in the least important principal component of the population data, removing them one at a time until the computation of $\theta(F^{\star})$ is numerically stable. The resulting parameter dimension, $p\coloneq\dim(\theta)$, and the features used in the functional are listed in Table 3 and Table 4, respectively.
We then standardise all continuous response and features using the population mean and variance, and one-hot encode all categorical features. Using population moments for standardisation ensures that $\theta$ remains on a common scale across repetitions.
E.3 Data setups
Synthetic data. The 11 synthetic setups are adapted from Wu and Martin (2023, Section 5): the features $x_{i}$ are i.i.d. and uniformly sampled from the hypercube $1,1]^{10}$. The true coefficients $\beta^{\star}$ are sampled once from $2,3]^{10}$ and held fixed across the 100 repetition of $z_{1:n}$. The response $y\mid x$ is generated as follows:
  • Continuous i.i.d. response (4 setups): We generate $n=20$ samples from $y_{i}=x^{\top}_{i}\beta^{\star}+\epsilon_{i},\,i=1,\ldots,n$. The errors $\epsilon_{i}$ are either i.i.d. ${\mathcal{N}}(0,1)$ or Student-$t$ with 5, 4, or 3 degrees of freedom. Note that $\theta(F^{\star})$ coincides with $(0,{\beta^{\star}}^{\top})$ in the case of i.i.d. $N(0,1)$ error.
  • Continuous response with heteroscedastic errors (3 setups): We generate $n=20$ samples from $y_{i}=x^{\top}_{i}\beta^{\star}+\epsilon_{i},\,i=1,\ldots,n$, where $\epsilon_{i}\sim{\mathcal{N}}(0,\sigma_{i}^{2})$. The standard deviation of the error $\sigma_{i}$ is determined by the first coordinate of $x_{i}$:
    $\sigma_{i}=\begin{cases}s_{\mathrm{left}},&\text{if }x_{i1}<\hat{\xi}_{0.25},\\ s_{\mathrm{mid}},&\text{if }\hat{\xi}_{0.25}\leq x_{i1}\leq\hat{\xi}_{0.75},\\ 1,&\text{if }x_{i1}>\hat{\xi}_{0.75},\end{cases}$
    where $\hat{\xi}_{0.25}$ and $\hat{\xi}_{0.75}$ are the empirical quartiles of $x_{11},\ldots,x_{n1}$. The hyperparameters $(s_{\text{left}},s_{\text{mid}})$ are arranged by their deviation from $\sigma=1$, ranging from weak to strong: $s_{1}=(0.25,0.5)$, $s_{2}=(0.05,0.25)$, and $s_{3}=(0.01,0.1)$.
  • Binary response (4 setups): We generate $n=100$ samples from $y_{i}\mid x_{i}\overset{\mathrm{iid}}{\sim}\mathrm{Bernoulli}(L(x^{\top}_{i}\beta^{\star})),\,i=1,\ldots,n$, with different link functions $L$. The link is either the logistic function:
    $L(u)=(1+\exp(-u))^{-1},$
    or the distribution function of a Gaussian-mixture model (GMM):
    $L(u)=0.7\,\Phi(u\mid a,1)+0.3\,\Phi(u\mid 2,1),$
    where $\Phi(\cdot\mid m,1)$ denotes the distribution function of a Gaussian with mean $m$ and unit variance, and $a\in\{0,-1,-2\}$. Note that $\theta(F^{\star})$ coincides with $(0,{\beta^{\star}}^{\top})$ in the case of logistic link.
Real data. We apply train_test_split in scikit-learn (Pedregosa et al., 2011) on the whole dataset, and use the training split as the dataset $z_{1:n}$. We repeat this procedure 100 times to generate repetitions of $z_{1:n}$. For datasets with highly skewed responses or features, we apply stratified sampling on these variables. For yeast and wine, which have highly skewed multi-class categorical responses, we also remove classes with fewer than 3% of observations. Multinomial logistic regression on a small dataset with an extremely skewed response remains difficult with our method, as the design matrix tends to be ill-conditioned when applying the Bayesian bootstrap on $x$.
The training sets are split from the population set with stratification on all categorical and numerical variables with very low unique values. In certain datasets with heavily skewed numerical variables, we also stratify according to the ‘bins’ of these numerical variables. The bin boundaries are determined to ensure that each bin contains approximately the same number of data points. The number of bins depends on the datasets and is chosen such that the marginal distribution of the training set roughly matches that of the population set.
The details of the 19 real-world datasets are provided in Tables 3 and 4.
Table 3: Details of the real-world datasets.
$z_{1:n}$ Data ID $n$ $p$ $n/p$ Population size Num. classes Num. cont. features Num. cat. features
concrete OpenML 44959 100 7 14.3 1030 0 8 0
quake OpenML 550 50 4 12.5 2178 0 3 0
airfoil OpenML 44957 50 5 10.0 1503 0 5 0
energy OpenML 44960 50 6 8.3 768 0 8 0
fish OpenML 44970 50 6 8.3 908 0 6 0
kin8nm OpenML 44980 50 9 5.6 8192 0 8 0
auction OpenML 44958 50 11 4.5 2043 0 5 2
grid OpenML 44973 50 12 4.2 10000 0 12 0
abalone OpenML 45042 20 5 4.0 4177 0 7 1
rice UCI 545 100 4 25.0 3810 2 7 0
sepsis OpenML 827 100 4 25.0 110341 2 2 1
banknote OpenML 1462 100 4 25.0 1372 2 4 0
mozilla OpenML 1046 100 5 20.0 15545 2 5 0
skin OpenML 1502 50 3 16.7 245057 2 3 0
blood OpenML 1464 50 4 12.5 748 2 4 0
phoneme OpenML 1489 50 6 8.3 5404 2 5 0
telescope UCI 159 50 9 5.6 19020 2 10 0
yeast OpenML 181 200 28 7.1 1350 5 8 0
wine OpenML 40498 200 40 5.0 4873 5 11 0
Table 4: Target name and the features included in the functional of interest.
$z_{1:n}$ Target Name Dropped features in the functional
concrete strength blast_furnace_slag, water
quake col_4
airfoil sound_pressure angle_of_attack
energy heating_load relative_compactness, surface_area, roof_area
fish LC50 MLOGP
kin8nm y
auction verification.time property.winner
grid stab p1
abalone Classnumberofrings Length, Diameter, Whole_weight, Viscera_weight, Shell_weight
rice Class Area, Perimeter, Major_Axis_Length, Convex_Area
sepsis hospital_outcome_1alive_0dead
banknote Class V3
mozilla state start
skin Class V2
blood Class V3
phoneme Class
telescope class fSize, fConc
yeast class_protein_localization erl, pox
wine Class V7, V8
E.4 Constructions of credible sets
Joint credible set Let $\widehat{\mu}$ and $\widehat{\Sigma}$ denote the empirical mean and covariance estimated from the posterior draws $\{\theta^{(l)}\}_{l=1}^{L}$. For martingale posteriors, we use a diagonal covariance structure $\widehat{\Sigma}=\operatorname{diag}(\widehat{\sigma}^{2}_{1},\ldots,\widehat{\sigma}^{2}_{p})$ due to the limited number of posterior samples available for estimating a full covariance matrix. Our credible set is then:
$\widehat{C}_{1-\alpha}=\Bigl\{\theta:(\theta-\widehat{\mu})^{\top}\widehat{\Sigma}^{-1}(\theta-\widehat{\mu})\leq\widehat{r}^{2}_{1-\alpha}\Bigr\},$
where
$\widehat{r}^{2}_{1-\alpha}=\operatorname{quantile}_{1-\alpha}\{(\theta^{(l)}-\widehat{\mu})^{\top}\widehat{\Sigma}^{-1}(\theta^{(l)}-\widehat{\mu})\}_{l=1}^{L},$
is an empirical cutoff. This empirical-cutoff ellipsoid provides a data-driven approximation of the high-posterior density set without relying on $\chi^{2}_{p,1-\alpha}$, as stated in (12).
Marginal credible interval. Let $\theta_{j}$ be the $j$-th dimension of $\theta$. The $(1-\alpha)$ credible interval is computed from the empirical quantiles of its posterior draws:
$\left[\operatorname{quantile}_{\alpha/2}\{\theta_{j}^{(l)}\}_{l=1}^{L},\;\;\operatorname{quantile}_{1-\alpha/2}\{\theta_{j}^{(l)}\}_{l=1}^{L}\right].$ (13)
Appendix F Diagnostics
Tracking each dimension of $\theta(F_{N})$ as $N$ grows is a common approach for diagnosing the convergence of $\theta(F_{N})$. However, this becomes impractical when $\theta$ is high-dimensional. Instead, we monitor convergence using the (scaled) expected $L_{1}$-norm between $\theta(F_{n})$ and $\theta(F_{N})$:
$\mathbb{E}_{F_{N}}\left[\frac{1}{p}\lVert\theta(F_{n})-\theta(F_{N})\rVert_{1}\right],$
where $p\coloneq\dim(\theta)$ and $\theta(F_{n})$ is the empirical risk minimiser (or maximum likelihood estimate of $z_{1:n}$ in our experiment). In practice, we simulate many realisations of $F_{N}$ to approximate this expectation. This expectation is shown in Figure 3 and corresponds to the black solid trajectories in Figure 6. In addition to the expected value, we also present the $L_{1}$-norm for each realisation of $F_{N}$ in Figure 6. While the paths corresponding to individual realisations of $F_{N}$ exhibit variability, they generally stabilise as $N$ increases, providing further empirical evidence that $\theta(F_{N})$ converges $\mathbb{P}$}-a.s..
For completeness, we also provide trace plots for each dimension of $\theta$ in the supplementary materials.
Figure 6: $L_{1}$-norm between $\theta(F_{n})$ and $\theta(F_{N})$ as $N$ increases. Each of the 30 plots corresponds to a realisation $z_{1:n}$ of a particular setup. Each of the grey lines corresponds to a realisation of $F_{N}$, and the black solid lines are their average.
Appendix G Marginal density and credible intervals of the posteriors
We present density plots of all posterior constructions for a single repetition of $z_{1:n}$. $95\%$ marginal credible intervals are shown as coloured horizontal bars in each plot.
Due to space constraints, we display only the following setups: setups:
  1. 1.
    Linear regression with synthetic Gaussian i.i.d. data (Figure 7);
  2. 2.
    Logistic regression with synthetic Bernoulli data with a logistic link (Figure 8);
  3. 3.
    Linear regression with the ‘abalone’ dataset (Figure 9);
  4. 4.
    Logistic regression with the ‘telescope’ dataset (Figure 10);
  5. 5.
    Multinomial logistic regression with the ‘yeast’ dataset (Figure 11).
The density plots for the remaining setups are provided in the supplementary material.
Figure 7: Marginal density plots of linear regression with the synthetic Gaussian i.i.d. data. For each posterior density, the 95% marginal credible interval is shown as a horizontal bar, and the posterior mean is marked. The solid and dashed black vertical lines correspond to $\theta(F^{\star})$ and $\theta(F_{n})$ respectively.
Figure 8: Marginal density plots of linear regression with the synthetic Bernoulli data with a logistic link. For each posterior density, the 95% marginal credible interval is shown as a horizontal bar, and the posterior mean is marked. The solid and dashed black vertical lines correspond to $\theta(F^{\star})$ and $\theta(F_{n})$ respectively.
Figure 9: Marginal density plots for the ‘abalone’ dataset. For each posterior density, the 95% marginal credible interval is shown as a horizontal bar, and the posterior mean is marked. The solid and dashed black vertical lines correspond to $\theta(F^{\star})$ and $\theta(F_{n})$ respectively.
Figure 10: Marginal density plots for the ‘telescope’ dataset. For each posterior density, the 95% marginal credible interval is shown as a horizontal bar, and the posterior mean is marked. The solid and dashed black vertical lines correspond to $\theta(F^{\star})$ and $\theta(F_{n})$ respectively.
Figure 11: Marginal density plots for the ‘yeast’ dataset. For each posterior density, the 95% marginal credible interval is shown as a horizontal bar, and the posterior mean is marked. The solid and dashed black vertical lines correspond to $\theta(F^{\star})$ and $\theta(F_{n})$ respectively.
Appendix H Coverage rate of the marginal credible interval
In addition to the joint credible sets discussed in the main text, we evaluate the performance of the $95\%$ marginal credible intervals (13) for each parameter. Figure 12 shows boxplots of the marginal coverage rates for each method across all parameters in our experiments. While most methods’ coverage is centred near the nominal 0.95 level, BB exhibits substantial under-coverage. Bayes also tends to undercover in logistic regression.
(a) Synthetic data
(b) Real-data (linear regression)
(c) Real-data (logistic regression)
Figure 12: Coverage of the 95% marginal credible intervals. Each data point in the boxplots represents the coverage for a single dimension of $\theta$ over 100 realisations of $z_{1:n}$.
The subsequent tables (Table 5 to Table 12) shows the empirical coverage and the median width for each parameter, computed over 100 data realisations. The method achieving at least the target coverage with the smallest size is highlighted in bold. If none of the methods achieve the target coverage, the method with the highest coverage is highlighted in bold instead. Summarising across all 30 setups and 293 individual parameters, TabMGP is bolded 152 times, followed by Copula (100), the classical Bayesian posterior (37) and Bayesian bootstrap (4). TabMGP performs particularly well in multinomial logistic regression relative to the competing methods.
Table 5: Coverage (rate) and median width (size) over 100 repetitions of marginal credible intervals for linear regression with synthetic i.i.d. data. The method achieving at least the target coverage with the smallest size is shown in bold. If none of the methods achieve the target coverage, the method with the highest coverage is shown in bold. The target coverage is 0.95.
TabMGP $B$ Copula Bayes
Setup Feature Name Rate Size Rate Size Rate Size Rate Size
intercept 0.97 0.73 0.83 0.30 1.00 0.63 1.00 0.87
x1 1.00 0.64 0.77 0.31 1.00 0.67 1.00 0.94
x10 0.97 0.78 0.79 0.28 0.98 0.62 1.00 0.90
x2 0.94 0.65 0.75 0.30 0.99 0.65 1.00 0.88
x3 0.99 0.65 0.76 0.32 0.99 0.65 1.00 0.90
x4 0.99 0.63 0.81 0.29 0.99 0.64 1.00 0.90
x5 0.99 0.65 0.76 0.31 1.00 0.67 1.00 0.91
x6 0.95 0.68 0.72 0.31 0.97 0.65 1.00 0.92
x7 0.96 0.68 0.67 0.30 0.98 0.65 1.00 0.86
x8 0.94 0.69 0.78 0.30 0.96 0.66 1.00 0.90
$N(0,1)$ x9 0.97 0.70 0.79 0.33 0.98 0.67 1.00 0.90
intercept 0.97 0.78 0.79 0.34 1.00 0.65 1.00 0.87
x1 1.00 0.67 0.78 0.36 0.99 0.67 1.00 0.95
x10 0.92 0.84 0.76 0.37 0.88 0.63 1.00 0.90
x2 0.96 0.70 0.76 0.37 0.96 0.67 0.99 0.89
x3 0.99 0.66 0.76 0.36 1.00 0.65 1.00 0.91
x4 1.00 0.66 0.77 0.34 0.97 0.65 1.00 0.91
x5 0.99 0.66 0.75 0.35 0.99 0.68 1.00 0.91
x6 0.98 0.74 0.76 0.36 0.97 0.66 1.00 0.92
x7 0.99 0.69 0.78 0.36 1.00 0.66 1.00 0.87
x8 0.94 0.73 0.68 0.34 0.97 0.64 1.00 0.89
$t_{5}$ x9 0.97 0.74 0.81 0.34 0.98 0.66 1.00 0.90
intercept 0.97 0.80 0.81 0.35 1.00 0.64 1.00 0.87
x1 1.00 0.68 0.79 0.38 0.99 0.67 1.00 0.95
x10 0.93 0.83 0.75 0.40 0.87 0.63 0.99 0.90
x2 0.94 0.68 0.76 0.39 0.96 0.66 0.99 0.88
x3 0.99 0.67 0.78 0.38 1.00 0.65 1.00 0.90
x4 1.00 0.68 0.78 0.36 0.98 0.65 1.00 0.92
x5 0.99 0.67 0.76 0.36 0.99 0.66 1.00 0.90
x6 0.99 0.75 0.75 0.38 0.97 0.67 1.00 0.91
x7 0.98 0.71 0.78 0.38 0.98 0.65 0.99 0.87
x8 0.94 0.71 0.70 0.34 0.97 0.64 0.99 0.89
$t_{4}$ x9 0.97 0.72 0.81 0.36 0.98 0.67 1.00 0.89
intercept 0.97 0.80 0.82 0.38 0.96 0.62 0.99 0.87
x1 1.00 0.68 0.79 0.40 0.99 0.66 0.99 0.95
x10 0.93 0.83 0.74 0.41 0.85 0.62 0.98 0.91
x2 0.95 0.67 0.76 0.42 0.96 0.66 0.97 0.89
x3 0.99 0.67 0.78 0.40 0.98 0.65 0.98 0.90
x4 1.00 0.68 0.78 0.38 0.98 0.65 0.98 0.90
x5 0.99 0.66 0.73 0.39 0.98 0.66 1.00 0.91
x6 0.98 0.73 0.73 0.41 0.94 0.65 0.96 0.92
x7 0.98 0.69 0.77 0.40 0.98 0.64 0.98 0.87
x8 0.92 0.71 0.71 0.37 0.93 0.62 0.97 0.90
$t_{3}$ x9 0.98 0.73 0.79 0.38 0.97 0.66 0.97 0.90
Table 6: Coverage (rate) and median width (size) over 100 repetitions of marginal credible intervals for linear regression with synthetic, dependent-error data. The method achieving at least the target coverage with the smallest size is shown in bold. If none of the methods achieve the target coverage, the method with the highest coverage is shown in bold. The target coverage is 0.95.
TabMGP $B$ Copula Bayes
Setup Feature Name Rate Size Rate Size Rate Size Rate Size
intercept 0.99 0.66 0.79 0.18 1.00 0.63 1.00 0.86
x1 1.00 0.55 0.73 0.19 1.00 0.67 1.00 0.94
x10 0.99 0.71 0.78 0.19 0.98 0.64 1.00 0.91
x2 0.97 0.61 0.73 0.18 1.00 0.65 1.00 0.89
x3 1.00 0.59 0.80 0.20 1.00 0.66 1.00 0.90
x4 1.00 0.57 0.77 0.19 1.00 0.66 1.00 0.91
x5 1.00 0.59 0.77 0.20 1.00 0.66 1.00 0.91
x6 0.98 0.64 0.67 0.19 0.99 0.65 1.00 0.92
x7 0.99 0.59 0.65 0.19 1.00 0.64 1.00 0.87
x8 0.97 0.63 0.72 0.19 0.98 0.66 1.00 0.89
$s_{1}$ x9 0.99 0.62 0.78 0.20 1.00 0.66 1.00 0.89
intercept 1.00 0.65 0.82 0.14 1.00 0.63 1.00 0.86
x1 0.99 0.54 0.62 0.16 1.00 0.66 1.00 0.95
x10 0.99 0.70 0.75 0.15 0.96 0.64 1.00 0.91
x2 0.99 0.58 0.78 0.14 1.00 0.65 1.00 0.89
x3 1.00 0.55 0.72 0.15 1.00 0.65 1.00 0.90
x4 1.00 0.55 0.76 0.15 1.00 0.67 1.00 0.91
x5 0.99 0.58 0.76 0.16 1.00 0.67 1.00 0.91
x6 1.00 0.61 0.64 0.15 0.99 0.65 1.00 0.92
x7 0.99 0.57 0.75 0.15 1.00 0.63 1.00 0.87
x8 0.98 0.61 0.66 0.15 1.00 0.66 1.00 0.89
$s_{2}$ x9 1.00 0.60 0.77 0.16 1.00 0.65 1.00 0.90
intercept 1.00 0.63 0.82 0.13 1.00 0.64 1.00 0.87
x1 0.99 0.55 0.62 0.15 1.00 0.66 1.00 0.95
x10 0.99 0.70 0.74 0.14 0.97 0.64 1.00 0.90
x2 0.98 0.58 0.78 0.13 1.00 0.65 1.00 0.89
x3 0.99 0.56 0.68 0.15 1.00 0.66 1.00 0.91
x4 1.00 0.57 0.69 0.14 1.00 0.67 1.00 0.91
x5 1.00 0.58 0.75 0.15 1.00 0.67 1.00 0.90
x6 1.00 0.61 0.62 0.15 1.00 0.65 1.00 0.92
x7 1.00 0.56 0.76 0.14 1.00 0.63 1.00 0.87
x8 0.98 0.60 0.67 0.14 1.00 0.66 1.00 0.90
$s_{3}$ x9 1.00 0.61 0.74 0.15 1.00 0.66 1.00 0.89
Table 7: Coverage (rate) and median width (size) over 100 repetitions of marginal credible intervals for logistic regression with synthetic data. The method achieving at least the target coverage with the smallest size is shown in bold. If none of the methods achieve the target coverage, the method with the highest coverage is shown in bold. The target coverage is 0.95.
TabMGP $B$ Copula Bayes
Setup Feature Name Rate Size Rate Size Rate Size Rate Size
intercept_1.0 0.96 1.11 0.88 1.36 0.98 1.67 0.84 0.93
x10_1.0 0.86 1.76 0.66 2.07 0.99 2.24 0.69 1.25
x1_1.0 0.97 0.97 0.86 1.44 0.99 1.69 0.77 0.95
x2_1.0 0.85 1.19 0.76 1.56 0.98 1.82 0.71 1.01
x3_1.0 0.95 1.04 0.81 1.41 1.00 1.69 0.80 0.97
x4_1.0 0.98 0.99 0.85 1.41 0.99 1.71 0.73 0.94
x5_1.0 0.94 1.04 0.86 1.44 0.99 1.73 0.81 0.95
x6_1.0 0.90 1.34 0.77 1.56 1.00 1.88 0.75 1.04
x7_1.0 0.89 1.21 0.79 1.54 1.00 1.85 0.72 1.00
x8_1.0 0.85 1.48 0.75 1.74 0.99 2.01 0.70 1.11
Logistic x9_1.0 0.89 1.44 0.73 1.67 0.99 1.94 0.69 1.08
intercept_1.0 0.90 1.48 0.73 2.02 0.99 2.16 0.70 1.24
x10_1.0 0.91 2.53 0.52 3.37 1.00 3.00 0.56 1.87
x1_1.0 0.96 1.21 0.82 1.84 1.00 1.98 0.74 1.14
x2_1.0 0.86 1.61 0.70 2.09 0.97 2.16 0.71 1.27
x3_1.0 0.91 1.33 0.79 1.79 0.99 1.96 0.75 1.16
x4_1.0 0.97 1.26 0.79 1.77 0.99 2.00 0.73 1.16
x5_1.0 0.94 1.24 0.84 1.76 0.97 1.89 0.79 1.16
x6_1.0 0.91 1.86 0.68 2.37 1.00 2.37 0.67 1.35
x7_1.0 0.93 1.60 0.74 2.02 0.99 2.11 0.75 1.26
x8_1.0 0.90 2.07 0.62 2.76 0.99 2.51 0.63 1.50
GMM(0)(0) x9_1.0 0.95 1.92 0.64 2.54 1.00 2.33 0.66 1.43
intercept_1.0 0.96 1.10 0.89 1.36 0.98 1.67 0.80 0.94
x10_1.0 0.90 1.78 0.66 1.96 1.00 2.25 0.71 1.26
x1_1.0 0.96 0.96 0.82 1.46 1.00 1.68 0.76 0.96
x2_1.0 0.85 1.24 0.77 1.54 0.99 1.80 0.76 1.02
x3_1.0 0.91 1.06 0.83 1.43 0.99 1.68 0.78 0.97
x4_1.0 0.96 0.99 0.82 1.39 0.99 1.69 0.74 0.95
x5_1.0 0.93 1.02 0.88 1.42 1.00 1.71 0.79 0.95
x6_1.0 0.88 1.36 0.75 1.59 1.00 1.86 0.72 1.04
x7_1.0 0.86 1.18 0.78 1.55 1.00 1.82 0.74 0.98
x8_1.0 0.86 1.51 0.68 1.81 0.99 2.02 0.66 1.11
GMM(1)(-1) x9_1.0 0.94 1.46 0.73 1.65 1.00 1.92 0.66 1.07
intercept_1.0 0.92 1.05 0.84 1.35 0.98 1.67 0.79 0.89
x10_1.0 0.85 1.42 0.73 1.48 1.00 1.82 0.77 1.03
x1_1.0 0.97 0.81 0.89 1.22 0.97 1.56 0.86 0.85
x2_1.0 0.87 0.94 0.78 1.28 0.99 1.60 0.77 0.89
x3_1.0 0.94 0.85 0.88 1.24 0.97 1.58 0.78 0.86
x4_1.0 0.99 0.84 0.91 1.21 1.00 1.54 0.78 0.85
x5_1.0 0.90 0.88 0.82 1.26 1.00 1.53 0.78 0.86
x6_1.0 0.91 1.05 0.84 1.33 1.00 1.67 0.77 0.91
x7_1.0 0.83 0.97 0.84 1.33 1.00 1.61 0.77 0.87
x8_1.0 0.81 1.18 0.78 1.37 1.00 1.74 0.74 0.92
GMM(2)(-2) x9_1.0 0.83 1.08 0.79 1.33 1.00 1.65 0.71 0.91
Table 8: Coverage (rate) and median width (size) over 100 repetitions of marginal credible intervals for linear regression with real data. The method achieving at least the target coverage with the smallest size is shown in bold. If none of the methods achieve the target coverage, the method with the highest coverage is shown in bold. The target coverage is 0.95.
TabMGP $B$ Copula Bayes
Setup Feature Name Rate Size Rate Size Rate Size Rate Size
age 0.83 0.36 0.83 0.35 0.83 0.32 0.81 0.29
cement 0.90 0.32 0.91 0.30 0.98 0.50 0.92 0.34
coarse_aggregate 0.91 0.29 0.91 0.27 0.99 0.47 0.94 0.30
fine_aggregate 0.95 0.29 0.95 0.26 1.00 0.45 0.96 0.31
fly_ash 0.96 0.34 0.92 0.32 0.99 0.63 0.96 0.36
intercept 0.91 0.29 0.92 0.26 1.00 0.55 0.95 0.28
concrete superplasticizer 0.91 0.37 0.85 0.35 0.99 0.49 0.90 0.36
col_1 1.00 0.23 0.88 0.44 0.99 0.50 0.95 0.45
col_2 0.93 0.25 0.94 0.51 1.00 0.80 0.88 0.41
col_3 1.00 0.24 0.95 0.50 1.00 0.81 0.91 0.40
quake intercept 0.96 0.47 0.95 0.50 1.00 1.02 0.96 0.39
chord_length 0.96 0.44 0.91 0.38 0.98 0.55 0.96 0.42
displacement_thickness 0.93 0.45 0.87 0.40 0.96 0.59 0.94 0.42
free_stream_velocity 0.94 0.40 0.93 0.35 0.99 0.55 0.95 0.41
frequency 0.90 0.48 0.87 0.43 0.84 0.47 0.88 0.42
airfoil intercept 0.93 0.42 0.93 0.36 0.99 0.63 0.95 0.40
glazing_area 0.93 0.25 0.86 0.15 1.00 0.34 1.00 0.42
glazing_area_distribution 0.96 0.25 0.89 0.15 0.99 0.33 1.00 0.43
intercept 0.92 0.28 0.88 0.15 1.00 0.45 1.00 0.40
orientation 0.96 0.24 0.88 0.15 1.00 0.34 1.00 0.41
overall_height 0.97 0.26 0.89 0.15 1.00 0.49 1.00 0.42
energy wall_area 1.00 0.25 0.96 0.12 1.00 0.35 1.00 0.43
CIC0 0.89 0.46 0.86 0.43 0.86 0.49 0.90 0.43
GATS1i 0.92 0.41 0.92 0.34 0.96 0.46 0.97 0.42
NdsCH 0.95 0.45 0.81 0.36 0.95 0.45 0.93 0.45
NdssC 0.94 0.46 0.89 0.44 0.94 0.47 0.93 0.48
SM1_Dz 0.89 0.48 0.83 0.38 0.92 0.48 0.88 0.45
fish intercept 0.93 0.43 0.83 0.35 0.99 0.54 0.90 0.41
intercept 0.94 0.47 0.91 0.38 0.90 0.35 0.94 0.42
theta1 0.96 0.41 0.85 0.38 0.86 0.36 0.92 0.44
theta2 0.98 0.40 0.86 0.37 0.88 0.35 0.96 0.43
theta3 0.96 0.52 0.90 0.36 0.78 0.34 0.94 0.43
theta4 0.96 0.42 0.90 0.39 0.90 0.34 0.93 0.43
theta5 0.82 0.44 0.87 0.40 0.86 0.33 0.91 0.42
theta6 0.92 0.43 0.79 0.39 0.78 0.34 0.85 0.43
theta7 0.94 0.42 0.84 0.38 0.84 0.35 0.91 0.43
kin8nm theta8 0.96 0.41 0.84 0.37 0.87 0.35 0.92 0.42
Table 9: Coverage (rate) and median width (size) over 100 repetitions of marginal credible intervals for linear regression with real data. The method achieving at least the target coverage with the smallest size is shown in bold. If none of the methods achieve the target coverage, the method with the highest coverage is shown in bold. The target coverage is 0.95.
TabMGP $B$ Copula Bayes
Setup Feature Name Rate Size Rate Size Rate Size Rate Size
intercept 0.93 0.74 0.89 0.53 1.00 1.70 0.99 0.87
process.b1.capacity 0.98 0.50 0.94 0.40 1.00 1.04 0.99 0.51
process.b2.capacity 0.90 0.39 0.80 0.32 1.00 0.84 0.96 0.46
process.b3.capacity 0.92 0.33 0.94 0.26 0.98 0.66 1.00 0.48
process.b4.capacity 0.93 0.40 0.86 0.32 1.00 1.20 0.99 0.45
property.price 0.91 0.45 0.81 0.37 1.00 1.03 0.99 0.52
property.product_2 0.91 1.39 0.87 1.08 0.84 1.61 0.94 1.26
property.product_3 0.93 1.06 0.89 0.85 0.80 1.61 0.99 1.48
property.product_4 0.96 1.04 0.89 0.84 0.88 1.39 1.00 1.44
property.product_5 0.94 1.14 0.89 0.89 0.97 1.67 1.00 1.64
auction property.product_6 0.94 0.99 0.93 0.77 0.82 1.50 0.99 1.34
g1 0.92 0.39 0.82 0.30 0.86 0.34 0.97 0.45
g2 0.98 0.40 0.88 0.30 0.94 0.33 1.00 0.45
g3 0.96 0.39 0.86 0.30 0.89 0.33 0.97 0.45
g4 0.98 0.40 0.92 0.31 0.93 0.33 0.97 0.45
intercept 0.92 0.40 0.91 0.30 0.86 0.33 0.98 0.44
p2 0.97 0.36 0.90 0.29 0.96 0.34 0.98 0.45
p3 0.99 0.36 0.91 0.29 0.97 0.34 0.99 0.44
p4 0.99 0.38 0.87 0.29 0.97 0.34 0.98 0.44
tau1 0.91 0.39 0.81 0.29 0.91 0.32 0.96 0.44
tau2 0.94 0.39 0.89 0.29 0.89 0.33 0.99 0.44
tau3 0.91 0.38 0.86 0.29 0.90 0.33 0.96 0.44
grid tau4 0.94 0.40 0.89 0.30 0.91 0.33 0.98 0.44
Height 0.99 1.53 0.65 1.23 0.96 1.36 0.78 1.49
Sex_1 0.92 1.74 0.78 1.53 0.95 1.78 0.92 1.92
Sex_2 0.99 1.48 0.81 1.47 0.95 1.55 0.91 1.63
Shucked_weight 0.96 1.40 0.68 1.23 0.92 1.48 0.82 1.36
abalone intercept 0.94 1.23 0.76 1.01 0.99 1.76 0.91 1.23
Table 10: Coverage (rate) and median width (size) over 100 repetitions of marginal credible intervals for logistic regression with real data. The method achieving at least the target coverage with the smallest size is shown in bold. If none of the methods achieve the target coverage, the method with the highest coverage is shown in bold. The target coverage is 0.95.
TabMGP $B$ Copula Bayes
Setup Feature Name Rate Size Rate Size Rate Size Rate Size
Eccentricity_Osmancik 0.91 5.36 0.37 7.52 0.98 14.46 0.62 4.81
Extent_Osmancik 0.99 1.19 0.93 2.11 1.00 4.14 0.93 1.40
Minor_Axis_Length_Osmancik 0.93 3.91 0.49 5.55 0.97 11.15 0.65 3.38
rice intercept_Osmancik 0.89 2.65 0.30 3.42 0.97 6.20 0.39 2.28
age_years_1 0.85 1.72 0.96 2.67 1.00 4.88 0.93 2.10
episode_number_1 1.00 1.19 0.93 2.54 1.00 3.28 0.99 1.83
intercept_1 1.00 1.77 0.98 2.91 1.00 6.42 0.97 1.95
sepsis sex_0male_1female_1_1 1.00 1.73 0.99 3.59 1.00 7.84 1.00 2.52
V1_2 0.96 2.24 0.96 2.62 1.00 2.85 0.98 1.96
V2_2 0.97 2.23 0.98 2.49 0.97 2.44 0.94 1.71
V4_2 0.98 1.99 0.98 2.25 0.99 2.30 0.92 1.38
banknote intercept_2 1.00 1.33 1.00 1.44 1.00 1.45 1.00 1.01
end_1 0.98 0.99 0.97 1.10 1.00 2.08 0.99 0.74
event_1 1.00 0.78 1.00 0.92 1.00 2.05 1.00 0.68
id_1 0.99 0.77 1.00 0.91 1.00 1.83 0.95 0.66
intercept_1 0.94 1.44 0.77 1.76 1.00 2.40 0.75 1.04
mozilla size_1 0.93 3.96 0.74 4.69 0.99 4.41 0.57 2.72
V1_2 1.00 3.00 1.00 3.29 1.00 3.23 0.99 1.59
V3_2 0.99 2.91 0.99 3.18 1.00 3.37 0.98 1.76
skin intercept_2 1.00 1.70 1.00 1.85 1.00 2.11 1.00 1.76
V1_2 0.82 1.80 0.91 2.64 0.98 4.18 0.83 1.63
V2_2 0.91 2.30 0.74 3.51 0.99 4.98 0.68 2.19
V4_2 0.81 1.76 0.79 2.74 1.00 4.01 0.77 1.91
blood intercept_2 1.00 1.48 0.93 2.14 1.00 3.68 0.92 1.28
V1_2 0.96 1.93 0.85 2.05 1.00 3.58 0.86 1.72
V2_2 0.95 1.84 0.87 2.15 0.98 3.44 0.80 1.46
V3_2 0.90 1.87 0.80 2.34 0.96 3.70 0.71 1.43
V4_2 0.91 1.68 0.80 1.97 0.95 3.68 0.75 1.24
V5_2 0.89 1.42 0.83 1.90 0.94 3.30 0.73 1.11
phoneme intercept_2 0.97 1.73 0.83 1.94 1.00 3.37 0.89 1.39
fAlpha_h 0.87 2.48 0.79 2.96 0.93 8.26 0.71 1.54
fAsym_h 0.96 2.55 0.78 3.38 1.00 7.77 0.72 2.26
fConc1_h 0.95 2.59 0.76 3.47 0.99 9.79 0.65 1.97
fDist_h 0.93 1.99 0.81 2.66 0.98 6.56 0.72 1.60
fLength_h 0.87 4.56 0.69 5.94 0.98 13.39 0.62 3.57
fM3Long_h 0.95 2.44 0.87 3.45 0.99 7.24 0.80 2.22
fM3Trans_h 0.97 2.21 0.88 3.12 1.00 6.63 0.81 1.92
fWidth_h 0.95 4.38 0.70 5.57 0.98 11.94 0.67 3.62
telescope intercept_h 0.97 1.93 0.88 2.54 1.00 7.44 0.83 1.56
Table 11: Coverage (rate) and median width (size) over 100 repetitions of marginal credible intervals for multinomial logistic regression with the ‘yeast’ data. The method achieving at least the target coverage with the smallest size is shown in bold. If none of the methods achieve the target coverage, the method with the highest coverage is shown in bold. The target coverage is 0.95.
TabMGP $B$ Bayes
Setup Feature Name Rate Size Rate Size Rate Size
alm_ME2 0.93 3.43 0.47 3.86 0.45 2.58
alm_ME3 0.95 3.16 0.56 4.22 0.70 2.81
alm_MIT 0.85 1.20 0.89 1.49 0.71 1.03
alm_NUC 0.93 0.79 0.90 0.97 0.90 0.79
gvh_ME2 0.94 2.39 0.56 3.27 0.57 2.19
gvh_ME3 0.94 1.41 0.82 1.95 0.73 1.49
gvh_MIT 0.92 0.94 0.86 1.20 0.77 0.87
gvh_NUC 0.94 0.61 0.93 0.86 0.86 0.64
intercept_ME2 0.89 6.35 0.39 10.10 0.43 4.42
intercept_ME3 0.95 3.53 0.54 4.87 0.67 3.01
intercept_MIT 0.97 1.08 0.87 1.30 0.82 0.94
intercept_NUC 1.00 0.70 1.00 0.81 1.00 0.62
mcg_ME2 0.90 2.74 0.61 4.14 0.59 2.47
mcg_ME3 0.92 1.63 0.82 2.11 0.81 1.64
mcg_MIT 0.90 0.90 0.90 1.12 0.78 0.85
mcg_NUC 0.96 0.66 0.94 0.89 0.86 0.64
mit_ME2 0.93 2.48 0.57 2.93 0.62 2.62
mit_ME3 0.97 1.99 0.84 2.65 0.77 2.07
mit_MIT 0.87 0.92 0.84 1.07 0.80 0.75
mit_NUC 0.93 0.76 0.90 0.93 0.87 0.70
nuc_ME2 0.98 3.30 0.55 3.92 0.65 3.26
nuc_ME3 0.74 1.67 0.74 2.23 0.69 1.77
nuc_MIT 0.94 1.34 0.83 1.61 0.76 1.28
nuc_NUC 0.89 0.81 0.81 1.04 0.72 0.70
vac_ME2 0.99 2.06 0.63 2.62 0.71 2.31
vac_ME3 0.97 1.60 0.87 2.22 0.91 1.98
vac_MIT 0.98 0.76 0.90 1.04 0.85 0.78
yeast vac_NUC 0.95 0.53 0.89 0.75 0.86 0.57
Table 12: Coverage (rate) and median width (size) over 100 repetitions of marginal credible intervals for multinomial logistic regression with the ‘wine’ data. The method achieving at least the target coverage with the smallest size is shown in bold. If none of the methods achieve the target coverage, the method with the highest coverage is shown in bold. The target coverage is 0.95.
TabMGP $B$ Bayes
Setup Feature Name Rate Size Rate Size Rate Size
V10_3 0.97 1.25 0.75 2.30 0.73 1.65
V10_4 0.96 1.27 0.71 2.18 0.71 1.61
V10_5 0.96 1.34 0.74 2.25 0.70 1.69
V10_6 0.96 1.64 0.79 3.03 0.73 2.09
V11_3 0.93 1.94 0.77 2.72 0.76 2.09
V11_4 0.97 1.98 0.73 2.78 0.78 2.09
V11_5 0.93 2.16 0.71 2.93 0.78 2.21
V11_6 0.93 2.83 0.64 4.22 0.67 2.85
V1_3 0.97 1.21 0.72 2.22 0.74 1.50
V1_4 0.90 1.26 0.74 2.20 0.72 1.48
V1_5 0.88 1.35 0.74 2.33 0.68 1.59
V1_6 0.91 1.73 0.74 3.17 0.74 2.19
V2_3 0.90 1.12 0.79 1.97 0.74 1.29
V2_4 0.81 1.30 0.73 2.06 0.68 1.36
V2_5 0.82 1.45 0.69 2.23 0.70 1.51
V2_6 0.84 1.99 0.67 3.12 0.66 2.15
V3_3 0.96 1.14 0.76 2.16 0.65 1.49
V3_4 0.97 1.18 0.72 2.21 0.63 1.49
V3_5 0.99 1.37 0.80 2.33 0.76 1.66
V3_6 1.00 1.85 0.74 3.19 0.71 2.37
V4_3 0.98 1.49 0.70 2.60 0.74 1.94
V4_4 0.96 1.50 0.74 2.66 0.75 1.93
V4_5 0.93 1.65 0.72 2.69 0.75 2.04
V4_6 0.90 2.17 0.71 3.78 0.64 2.66
V5_3 0.98 1.33 0.68 2.46 0.74 1.74
V5_4 0.96 1.37 0.77 2.55 0.77 1.77
V5_5 0.94 1.98 0.82 2.90 0.81 2.22
V5_6 0.97 2.47 0.74 4.78 0.75 3.36
V6_3 0.79 1.77 0.69 3.10 0.56 1.75
V6_4 0.84 1.78 0.70 3.06 0.57 1.77
V6_5 0.84 1.92 0.72 3.19 0.60 1.85
V6_6 0.83 2.15 0.70 4.07 0.59 2.37
V9_3 0.97 1.29 0.75 2.50 0.64 1.72
V9_4 0.98 1.34 0.72 2.50 0.65 1.74
V9_5 0.95 1.46 0.71 2.57 0.65 1.82
V9_6 0.94 1.81 0.73 3.51 0.70 2.33
intercept_3 0.97 2.87 0.42 5.76 0.55 2.76
intercept_4 0.99 2.90 0.37 5.72 0.49 2.75
intercept_5 0.98 3.00 0.53 5.91 0.56 2.82
wine intercept_6 0.99 3.94 0.81 9.06 0.69 4.35